OpenStudy (anonymous):

Please help! Sara saved $1,000 in the bank. She decided to leave it there for 5 years. Find the amount of money Sara will have in her account after 5 years if she chooses 6.8% interest compounded daily. (365 times a year-leap year not included).

OpenStudy (anonymous):

The Compound Interest Equation P = C (1 + r/n) nt where P = future value C = initial deposit r = interest rate (expressed as a fraction: eg. 0.06) n = # of times per year interest in compounded t = number of years invested Source: http://math2.org/math/general/interest.htm

OpenStudy (anonymous):

The last nt should be a power, sorry. P = C (1 + r/n)^nt

OpenStudy (anonymous):

So, in your question, P=1000(1+0.068/365)^(365*5)

OpenStudy (anonymous):

And, my calculator just broke, so you'll have to figure it out on your own, lol.

OpenStudy (anonymous):

P=1000(1+0.068/365)^365(5)?